OpenStudy (anonymous):

. Mr. Piñeda used 540 pounds of mulch to cover his front flowerbed, which has an area of 72 square feet. How many pounds of mulch will he need to cover his back flowerbed, which has an area of 180 square feet? (1 point)7.5 pounds 216 pounds 1,350 pounds 1,180 pounds 4. Last week a certain brand of bottled water cost $1.25 for a 16-ounce bottle. This week the water is on sale for $1.00 for a 16-ounce bottle. What is the percent decrease in the price of this bottled water? (1 point)16% 20% 13% 25% 5. Mrs. Guzman bought 3 table lanps for $129.90

OpenStudy (kamille):

1st question: You can make a proportion: You know, that he used 540 punds to cover area of 72 So he need to use x pounds to cover area of 180 \[\frac{ 540 }{ 72 }=\frac{ x }{ 180 }\] Now you need to dinf the x

OpenStudy (kamille):

DO you know how to do it?

OpenStudy (anonymous):

no i do not

OpenStudy (kamille):

Do you understand why I made a proportion?

OpenStudy (kamille):

\[\frac{ 540 }{72 }=\frac{ x }{ 180 }|*180\] \[\frac{ 540 * 180 }{ 72 }=x\] x=1350